Kasia Manufacturing Company has developed the following standards for one of their products:STANDARD VARIABLE COST CARD, ONE UNIT OF PRODUCT

Direct materials:  20 square meters at RM5 per square foot RM100.00

Direct labor:       8 hours at RM6 per hour- 48.00

Variable overhead:  8 hours at RM3 per hour-24.00

Total standard variable cost per unit -RM172.00

The company records materials price variances at the time of purchase.

The following activities occurred during the month of July:

Materials purchased:150,000 square metersat RM5.25 per sq. foot

Materials used:96,000 square meters

Units produced:5,000 units

Direct labor:41,000 hours at RM6.55 per hour

Actual variable overhead:RM238,000

Required:

Compute the following variances and indicate whether they are favorable (F) or unfavorable (U).

a. Material price variance? b.Material usage variance? c.Direct labourrate variance? d.Direct labor efficiency variance? e.Variable overhead spending variance? f. Variable overhead efficiency variance?  


1
Expert's answer
2021-12-14T15:57:37-0500

a.

Price variance = (Actual unit cost - Standard unit cost) * Actual Quantity Purchased

Price variance = "(5.25-5)\cdot 150000=37500"


b.

usage variance = (standard quantity of material allowed for production – actual quantity used) × standard price per unit of material

usage variance = "(20\cdot5000-96000)\cdot5=20000"


c.

Direct labour rate variance = (standard hours allowed for production – actual hours taken) × standard rate per direct labour hour

Direct labour rate variance = "(8\cdot5000-41000)\cdot6=-6000"


d.

Direct labor efficiency variance = standard labor hours/actual amount of time worked

Direct labor efficiency variance = "8\cdot5000/41000=0.9756=97.56\%"


e.

Variable overhead spending variance = (Actual variable overhead rate - Standard variable overhead rate) x Actual hours worked during the period

Variable overhead spending variance = "238-24=214"


f)

Variable overhead efficiency variance = Standard variable overhead/Actual variable overhead

Variable overhead efficiency variance = "24/238 =0.1008=10.08\%"